[Brahms-l] Shift report 20050524 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 24 2005 - 23:56:48 EDT
Shifter leader was kh
Shifters were og


Beam during most of the shift.  At shift start, CC counters powered by BBL were in a tripped state.  We cycled the power on BB3 and then turned the voltages up and that seemed to be the end of the problem.
Achieved runs 14901 through 14911.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 24 23:57:00 2005

This archive was generated by hypermail 2.1.8 : Tue May 24 2005 - 23:57:12 EDT